### v2

# Measurement of $\Sigma^{+}$ transverse polarization in $e^{+}e^{-}$ collisions at $\sqrt{s}=3.68-3.71$ GeV

Abstract
Using $e^{+}e^{-}$ collision data collected with the BESIII detector at seven energy points ranging from 3.68 to 3.71 GeV and corresponding to an integrated luminosity of $652.1~{}{\rm pb^{-1}}$, we present an energy-dependent measurement of the transverse polarization, relative phase and modulus ratio of the electromagnetic form factors of the $\Sigma^{+}$ hyperon in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ reaction. These results are helpful to understand the production mechanism of the $\Sigma^{+}$-$\bar{\Sigma}^{-}$ pairs.
1 Introduction
As early as 1960, the significance of baryon structure was widely acknowledged Cabibbo:1961sz . However, at present, understanding the structure of baryons remains a great challenge. Apart from protons and neutrons, knowledge has been relatively scarce. The proton, being a stable particle, allows for the determination of the space-like electromagnetic form factors (EMFFs) via elastic electron-proton scattering Perdrisat:2006hj . More recently, $e^{+}e^{-}$collisions, resulting in the production of baryon pairs, provide an ideal experimental framework to study baryon structure. This is because the $e^{+}e^{-}$ process offers access to the time-like EMFFs via virtual photon production, thereby facilitating the quantitative assessment of baryonic electromagnetic structure. Experimentally accessible time-like EMFFs are intimately connected with more intuitive space-like quantities, such as charge and magnetization densities, through the dispersion relation Belushkin:2006qa ; Yan:2023yff . Baryon pairs with a spin of 1/2 produced by the $e^{+}e^{-}$ process via a virtual photon are elegantly described by two distinct EMFFs: the electric form factor $G_{E}$ and magnetic form factor $G_{M}$ Huang:2021xte ; Aubert:2007uf . These form factors are functions of the square of the four-momentum transfer, denoted as $q^{2}$. In the time-like region, where $q^{2}>0$, the EMFFs exhibit non-zero imaginary components. Consequently, when $G_{E}$ and $G_{M}$ are different, they give rise to a relative phase denoted as $\Delta\Phi=\Phi_{E}-\Phi_{M}$, where $\Phi_{E}$ and $\Phi_{M}$ are the phases of $G_{E}$ and $G_{M}$, respectively.
In accordance with the Phragmén-Lindelöf theorem titchmarsh1939theory , as $q^{2}\to\infty$, the asymptotic behavior of the time-like EMFFs can be obtained from their corresponding space-like counterparts. In the space-like region, the EMFFs are assumed to be real. Consequently, as $q^{2}\to\infty$, the time-like EMFFs must also be real, and the relative phase of the EMFFs should approach an integer multiple of $\pi$. In the case where $q^{2}$ does not approach the limit, the phase can be any value other than zero, which causes the final state to be polarized, even if the initial state is unpolarized Dubnickova:1992ii . This allows the determination of the EMFFs, as the polarization and cross-section are functions of form factors themselves Faldt:2017kgy ; Cao:2021asd ; Dai:2023vsw . Recent results from BESIII BESIII:2020uqk , BaBar Aubert:2007uf and Belle Belle:2022dvb , focusing solely on measuring the effective form factor, have reported consistent results. Additionally, the CLEO collaboration has made significant progress in determining the cross sections and the EMFFs of various baryon pairs ($p$, $\Lambda$, $\Sigma^{0}$, $\Sigma^{+}$, $\Xi^{0}$, $\Xi^{-}$, and $\Omega^{-}$) Dobbs:2014ifa ; Dobbs:2017 . Their conclusions regarding EMFFs and di-quark correlations Jafee:2005 ; Jafee:2003 rely on the assumption that one-photon exchange dominates the production process and that charmonia contributions are negligible. Furthermore, the BESIII collaboration has measured the cross sections for several baryon pairs ($\Lambda$, $\Sigma^{0}$, $\Xi^{-}$, $\Sigma^{\pm}$, and $\Xi^{0}$) near the production threshold BESIII:2017hyw ; BESIII:2020uqk ; BESIII:2020ktn ; BESIII:2021aer ; BESIII:2021rkn ; BESIII:2022kzc and above open charm threshold BESIII:2019cuv ; BESIII:2021ccp ; BESIII:2023rse ; BESIII:2024umc ; BESIII:2024ogz . However, experimental measurements of the relative phase between $G_{E}$ and $G_{M}$ are still relatively scarce.
For spin 1/2 baryon pairs produced by the $e^{+}e^{-}$ process via vector charmonia, where the above statements concerning EMFFs also apply, the formalism is described in ref. Faldt:2017kgy . In these cases, the amplitudes include the EM-psionic form factors, $G^{\Psi}_{E}$ and $G^{\Psi}_{M}$ BESIII:2021cvv ; BESIII:2023euh . While the EM-psionic form factors describe a different process, the form of the hadron current matrix element for the charmonia process is the same as that for the virtual photon one Faldt:2017kgy ; Faldt:EPJ . Previous studies often neglected the polarization effects of hyperons Ablikim:2016iym ; Ablikim:2016iym-01 ; Ablikim:2016iym-02 ; Ablikim:2016iym-03 ; Wang:2018kdh ; Wang:2021lfq . However, $\Sigma^{+}$ polarization was recently observed and measured in the $e^{+}e^{-}\to J/\psi$, $\psi(3686)\to\Sigma^{+}\bar{\Sigma}^{-}$ processes by the BESIII collaboration BESIII:2020fqg ; BESIII:2023sgt . The results not only reveal a non-zero relative phase but also demonstrate that the phase changes sign at the mass of the $\psi(3686)$ resonance compared to the value measured at the $J/\psi$ resonance. Subsequently, the polarization effects of $\Lambda$ and $\Xi^{-}$ were also observed and measured by the BESIII collaboration. The $\Lambda$ polarization was observed in the $e^{+}e^{-}\to\Lambda\bar{\Lambda}$ process at the $J/\psi$, $\psi(3770)$, and off-resonance regions BESIII:2021cvv ; BESIII:2023euh ; BESIII:2022yprl ; Ablikim:2018zay . Additionally, a non-zero polarization has been observed for $\Xi^{-}$ in the $e^{+}e^{-}\rightarrow J/\psi$, $\psi(3686)\rightarrow\Xi^{-}\bar{\Xi}^{+}$ processes BESIII:2021ypr ; BESIII:2022_xi_psip ; Liu:2023xhg . However, for $\Xi^{0}$, completely different effects were observed in processes $e^{+}e^{-}\rightarrow J/\psi\rightarrow\Xi^{0}\bar{\Xi}^{0}$ BESIII:2023drj and $e^{+}e^{-}\rightarrow\psi(3686)\rightarrow\Xi^{0}\bar{\Xi}^{0}$ BESIII:2023lkg . The data samples used in this analysis correspond to an integrated luminosity of 652.1 pb-1, collected at the center-of-mass (CM) energies of $\sqrt{s}=$ 3.682, 3.683, 3.684, 3.685, 3.687, 3.691, and 3.710 GeV with the BESIII detector Ablikim:2009aa in symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring BEPCII . The above energy points around $\psi(3686)$ resonance, which are used to study the energy dependence of $\Delta\Phi$, are particularly intriguing, as the production occurs through a combination of one-photon exchange BESIII:2019nep , mixing with the $\psi(3686)$ resonance BESIII:2021cvv , and resonance dominance Ablikim:2018zay ; BESIII:2022yprl . Since the energy range studied here includes both the virtual photon vector charmonium processes, $G_{E}^{\gamma/\Psi}$ and $G_{M}^{\gamma/\Psi}$ are used to represent the form factors in the following, and the methods used in the following apply to both processes.
2 BESIII detector and Monte Carlo simulation
[…]
To evaluate detection efficiencies and estimate backgrounds, simulated data samples are produced using geant4-based Monte Carlo (MC) software GEANT4 , which incorporates the geometric description of the BESIII detector Huang:2022wuo as well as the detector response. The simulation models the beam energy spread and initial-state radiation (ISR) effect in the $e^{+}e^{-}$ annihilation process with kkmc KKMC . The detection efficiency for the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ process is determined through MC simulations. For each of the 7 energy points ranging from 3.68 to 3.71 GeV, a sample of 100,000 events is simulated with a uniform phase space (PHSP) distribution. The $\Sigma^{+}(\bar{\Sigma}^{-})$ baryon and its subsequent decays are simulated using evtgen evtgen2 ; EVTGEN with a PHSP model.
3 Event selection
A double-tag technique is employed in the event selection involving the decays of $\Sigma^{+}\to p\pi^{0}$ and $\bar{\Sigma}^{-}\to\bar{p}\pi^{0}$ with the subsequent decay $\pi^{0}\rightarrow\gamma\gamma$. Hence, in the final state, there are two charged particles, a proton and an anti-proton, along with four photons utilized for the reconstruction of two $\pi^{0}$s. Consequently, suitable candidates must meet the following event selection criteria.
Charged tracks detected in the MDC are required to be within a polar angle ($\theta$) range of $|\!\cos\theta|<0.93$, where $\theta$ is defined with respect to the $z$-axis, which is the symmetry axis of the MDC. At least two oppositely charged tracks, which must be well reconstructed in the MDC with good helix fits, are required. There is no vertex requirement. Since (anti-)protons from $\Sigma^{+}(\bar{\Sigma}^{-})$ decays can be distinguished from other charged particles by requiring the momentum to be greater than $0.5~{}{\rm GeV}/c$, there is no additional particle identification (PID) requirement.
Showers deposited in the EMC are used to reconstruct $\pi^{0}$s. The deposited energy of each shower must be more than 25 MeV in the barrel region ($|\!\cos\theta|<0.80$) and more than 50 MeV in the end cap region ($0.920.86<|\!\cos\theta|<$ 0.92). To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns. After selection, at least four photons are required.
To select the correct combination of proton, anti-proton, and $\pi^{0}\pi^{0}$ candidates, a six-constraint (6C) kinematic fit is applied to all $p\bar{p}\gamma\gamma\gamma\gamma$ combinations in each event. The 6C kinematic fit conserves energy and momentum while constraining the invariant mass of photon combinations to the known $\pi^{0}$ mass PDG2020 . The $p\bar{p}\pi^{0}\pi^{0}$ combination with the smallest $\chi^{2}$ is retained, and $\chi_{\rm 6C}^{2}<100$ is required to suppress the background. This requirement is determined by analyzing the figure-of-merit (FOM) $N_{\rm sig}/\sqrt{N_{\rm sig}+N_{\rm bkg}}$, where $N_{\rm sig}$ and $N_{\rm bkg}$ represent the numbers of signal and background events, respectively, both based on MC simulation. To match $p$ and $\bar{p}$ with the correct $\pi^{0}$, the $\Sigma^{+}\bar{\Sigma}^{-}$ pair with the minimum difference $\sqrt{(M_{p\pi^{0}}-m_{\Sigma^{+}})^{2}+(M_{\bar{p}\pi^{0}}-m_{\bar{\Sigma}^{-}})^{2}}$ is selected for further analysis. Here, $M_{p\pi^{0}(\bar{p}\pi^{0})}$ is the invariant mass of the $p\pi^{0}(\bar{p}\pi^{0})$ combination, and $m_{\Sigma^{+}(\bar{\Sigma}^{-})}$ is the known $\Sigma^{+}$ mass PDG2020 . To eliminate the primary background originating from the $e^{+}e^{-}\rightarrow\pi^{0}\pi^{0}J/\psi\rightarrow p\bar{p}\pi^{0}\pi^{0}$ process, the veto requirement $|M_{\pi^{0}\pi^{0}}^{\rm Recoil}-m_{J/\psi}|>15$ MeV/$c^{2}$ is applied. Here, $M_{\pi^{0}\pi^{0}}^{\rm Recoil}$ is the recoil mass of the $\pi^{0}\pi^{0}$ combination and $m_{J/\psi}$ is the mass of $J/\psi$ PDG2020 . This requirement is also determined by FOM optimization. Figure 1 shows the distribution of $M_{\bar{p}\pi^{0}}$ versus $M_{p\pi^{0}}$ of the accepted candidates summed over all energy points. The $M_{p\pi^{0}(\bar{p}\pi^{0})}$ candidates are required to be within the range of $[m_{\Sigma^{+}}-4\sigma,m_{\Sigma^{+}}+3\sigma]$ MeV/$c^{2}$, denoted by $S$ in figure 1, where the resolution $\sigma$ is determined by a one-dimensional fit with the Crystal-Ball function Oreglia:1980cs . Due to the energy leakage of the photon in the EMC, the signal shape is asymmetric.
Figure 1: Two-dimensional distribution of $M_{\bar{p}\pi^{0}}$ versus $M_{p\pi^{0}}$ of the accepted candidates summed over all energy points, where the red square marked with $S$ indicates the signal region, and the pink squares marked with $B_{i}~{}(i=1,2,3,4)$ show the selected background regions.
After applying the event selection criteria to the data, the remaining background mainly comes from non-$\Sigma^{+}(\bar{\Sigma}^{-})$ events, such as the non-resonant process $e^{+}e^{-}\rightarrow\pi^{0}\pi^{0}p\bar{p}$. The number of background events is estimated using the sideband method, i.e. $\sum^{4}_{i=1}B_{i}/4$ for the $M_{p\pi^{0}}$ and $M_{\bar{p}\pi^{0}}$ windows, where $i$ runs over the four regions shown in figure 1, and the exact ranges are defined as $B_{1}$: $M_{p\pi^{0}}~{}\in$ [1.119, 1.154] GeV/$c^{2}$ and $M_{\bar{p}\pi^{0}}~{}\in$ [1.219, 1.254] GeV/$c^{2}$, $B_{2}$: $M_{p\pi^{0}}~{}\in$ [1.219, 1.254] GeV/$c^{2}$ and $M_{\bar{p}\pi^{0}}~{}\in$ [1.219, 1.254] GeV/$c^{2}$, $B_{3}$: $M_{p\pi^{0}}~{}\in$ [1.119, 1.154] GeV/$c^{2}$ and $M_{\bar{p}\pi^{0}}~{}\in$ [1.119, 1.154] GeV/$c^{2}$, $B_{4}$: $M_{p\pi^{0}}~{}\in$ [1.219, 1.254] GeV/$c^{2}$ and $M_{\bar{p}\pi^{0}}~{}\in$ [1.119, 1.154] GeV/$c^{2}$.
The signal yield combining all energy points is $898\pm 30$ (stat.) events. The numbers of signal events for each energy point are listed in table 1. The number of background events estimated with the sideband method is $36\pm 3$ (stat.), which is a background level of 3.85%.
4 $\Sigma^{+}$ polarization
The exclusive process $e^{+}e^{-}\rightarrow\gamma^{*}/\Psi\rightarrow\Sigma^{+}\bar{\Sigma}^{-}\rightarrow p\bar{p}\pi^{0}\pi^{0}$ can be fully described by the $\Sigma^{+}$ scattering angle, $\theta_{\Sigma^{+}}$, in the CM system of the $e^{+}e^{-}$ reaction and the $p$ and  $\bar{p}$ directions, $\bm{\hat{n}_{1}}$ and $\bm{\hat{n}_{2}}$, respectively, in the rest frames of their parent particles. Here $\gamma^{*}/\Psi$ indicates that the process $e^{+}e^{-}\rightarrow\Sigma^{+}\bar{\Sigma}^{-}$ occurs via a pure EM process or a $\psi$ resonance. The components of these vectors are expressed using a coordinate system ($x_{\Sigma^{+}},y_{\Sigma^{+}},z_{\Sigma^{+}}$) as shown in figure 2. A right-handed system for each hyperon decay is defined with the $z$ axis along the $\Sigma^{+}$ momentum $_{\Sigma^{+}}=-\textbf{p}_{\bar{\Sigma}^{-}}=\textbf{p}$ p in the CM system. The $y$ axis is taken as the normal to the scattering plane, $\textbf{k}_{e^{-}}\times\textbf{p}_{\Sigma^{+}}$, where $_{e^{-}}=-\textbf{k}_{e^{+}}=\textbf{k}$ k is the electron beam momentum in the CM system.
Figure 2: Definition of the coordinate system describing the $e^{+}e^{-}\rightarrow\gamma^{*}/\Psi\rightarrow\Sigma^{+}\bar{\Sigma}^{-}\rightarrow p\bar{p}\pi^{0}\pi^{0}$ reaction. The $\Sigma^{+}$ particle is emitted along the $z_{\Sigma^{+}}$ axis direction, and the $\bar{\Sigma}^{-}$ in the opposite direction. The $y_{\Sigma^{+}}$ axis is perpendicular to the plane of $\Sigma^{+}$ and $e^{-}$, and the $x_{\Sigma^{+}}$ axis is defined by a right-handed coordinate system. The $\Sigma^{+}$ decay product, the proton, is measured in this coordinate system.
The joint decay angular distribution of the process $e^{+}e^{-}\rightarrow\gamma^{*}/\Psi\rightarrow\Sigma^{+}\bar{\Sigma}^{-}\rightarrow p\bar{p}\pi^{0}\pi^{0}$, involving spin-entangled $\Sigma^{+}\bar{\Sigma}^{-}$, is expressed as Faldt:2017kgy
${\cal{W}}(\bm{\xi};\bm{\Omega})=$ $\mathcal{F}_{0}(\xi)+\eta\mathcal{F}_{5}(\xi)+\alpha_{\Sigma^{+}}\alpha_{\bar{\Sigma}^{-}}$
$\times$ $[\mathcal{F}_{1}(\xi)+\sqrt{1-\eta^{2}}\cos(\Delta\Phi)\mathcal{F}_{2}(\xi)+\eta\mathcal{F}_{6}(\xi)]$
$+$ $\sqrt{1-\eta^{2}}\sin(\Delta\Phi)[\alpha_{\Sigma^{+}}\mathcal{F}_{3}(\xi)+\alpha_{\bar{\Sigma}^{-}}\mathcal{F}_{4}(\xi)],$ (1)
[…]
${{F}}_{1}=$ $\sum_{i=1}^{N_{k}}(\sin^{2}\!\theta_{\Sigma^{+}}n^{i}_{1,x}n^{i}_{2,x}+\cos^{2}\!\theta_{\Sigma^{+}}n^{i}_{1,z}n^{i}_{2,z}),$
${{F}}_{2}=$ $\sum_{i=1}^{N_{k}}\sin\!\theta_{\Sigma^{+}}\cos\!\theta_{\Sigma^{+}}(n^{i}_{1,x}n^{i}_{2,z}+n^{i}_{1,z}n^{i}_{2,x}),$
${{F}}_{3}=$ $\sum_{i=1}^{N_{k}}\sin\!\theta_{\Sigma^{+}}\cos\!\theta_{\Sigma^{+}}n^{i}_{1,y},$ (2)
${{F}}_{4}=$ $\sum_{i=1}^{N_{k}}\sin\!\theta_{\Sigma^{+}}\cos\!\theta_{\Sigma^{+}}n^{i}_{2,y},$
${{F}}_{5}=$ $\sum_{i=1}^{N_{k}}(n^{i}_{1,z}n^{i}_{2,z}-\sin^{2}\!\theta_{\Sigma^{+}}n^{i}_{1,y}n^{i}_{2,y}),$
where, $N_{k}$ is the number of events in the $k^{\rm th}$ $\cos\theta_{\Sigma^{+}}$ interval, $n_{1,j}~{}(n_{2,j})~{}(j=x,y,z)$ represents the component of vector $\mathbb{\hat{n}}_{1}$ ($\mathbb{\hat{n}}_{2}$) in the coordinate system ($x_{\Sigma^{+}}$, $y_{\Sigma^{+}}$, $z_{\Sigma^{+}}$), and $i$ is the index from 1 to $N_{k}$.
The modulus ratio of the form factors, $R$ BESIII:2021cvv , can be described by the angular distribution parameter $\eta$.
$R=\frac{|G_{E}^{\gamma/\Psi}|}{|G_{M}^{\gamma/\Psi}|}=\sqrt{\frac{\tau(1-\eta)}{1+\eta}},$ (3)
where $\tau=\frac{s}{4m_{\Sigma^{+}}^{2}}$ and $s$ is the square of the CM energy. If the initial state is unpolarized, and the production process is either strong or electromagnetic and hence parity-conserving, then a non-zero polarization is only possible in the transverse direction $y$. The polarization is given by Faldt:2017kgy
$P_{y}=\frac{\sqrt{1-\eta^{2}}\sin\theta_{\Sigma^{+}}\cos\theta_{\Sigma^{+}}}{1+\eta\cos^{2}\theta_{\Sigma^{+}}}\sin(\Delta\Phi).$ (4)
To determine $\eta$ and $\Delta\Phi$, an unbinned maximum likelihood fit is performed, where the decay parameters $\alpha_{\Sigma^{+}}$ and $\alpha_{\bar{\Sigma}^{-}}$ are fixed to the values $-0.994$ and 0.994, respectively, obtained from the average in ref. BESIII:2020fqg , assuming $CP$ conservation. In the fit, the likelihood function $\mathscr{L}$ is constructed from the probability function, ${\cal{P}}({\bm{\xi}}_{i})$, for event $i$ characterized by the measured angles $\bm{\xi}_{i}$
$\mathscr{L}=\prod_{i=1}^{N}{\cal{P}}({\bm{\xi}}_{i},{\bm{\Omega}})=\prod_{i=1}^{N}{\cal{C}}{\cal{W}}({\bm{\xi}}_{i},{\bm{\Omega}})\epsilon(\bm{\xi}_{i}),$ (5)
where $N$ is the number of events in the signal region, and $\epsilon(\bm{\xi}_{i})$ is the detection efficiency. For the ISR effect at the higher energy points 3.691 and 3.710 GeV, MC studies are performed where the input cross section for $e^{+}e^{-}\rightarrow\Sigma^{+}\bar{\Sigma}^{-}$ for calculating the ISR effect is taken from ref. borncs . The ISR effect at these two energy points brings absolute differences of 0.02 and 0.04 rad for $\eta$ and $\Delta\Phi$, respectively, which are negligible. The normalization factor $\mathcal{C}=\frac{1}{N_{\mathrm{MC}}}\sum_{j=1}^{N_{\mathrm{MC}}}{\cal{W}}({\bm{\xi}}^{j},{\bm{\Omega}})$ is given by the sum of the corresponding angular distribution function $\cal{W}$ using the accepted MC events, $N_{\mathrm{MC}}$, and the difference in efficiency between data and MC simulations is taken into account as a systematic uncertainty, described later. The minimization of the function
$\mathscr{S}=-\mathrm{ln}\mathscr{L}_{\rm data}$ (6)
is performed with RooFit Roofit . Here, $\mathscr{L}_{\rm data}$ is the likelihood function of events selected in the signal region.
Figure 3 shows the distributions of the five ${F}_{k}(k=1,2,...,5)$ moments Faldt:2017kgy with respect to $\cos\theta_{\Sigma^{+}}$, divided into 10 intervals, and the weighted PHSP MC results corrected by the global fit. The numerical fit results and the weighted average values are summarized in table 1. Here, the weighted average is calculated by $\frac{1}{w}\sum_{i=1}^{n}w_{i}x_{i}$ PDG2020 , where $n$ is the number of energy points, $w_{i}=1/\sigma_{i}^{2}$, $x_{i}$ and $\sigma_{i}$ are the measured values and their uncertainties at each energy point, respectively, and $w=\sum_{i}w_{i}$. The relative phases of the form factors at each energy point are measured to be different from zero with significances of less than $3\sigma$. The significance is estimated by comparing the likelihoods of the baseline fit and the one defined assuming no polarization p:value . The calculation of significance also accounts for the systematic uncertainties in the decay parameter, mass window, and background. After evaluating the significance following each systematic uncertainty, we select the smallest value as a conservative estimation. The significance of each energy point is also summarized in table 1.
Figure 3: Distributions of ${F}_{k}(k=1,2,...,5)$ moments with respect to $\cos\theta_{\Sigma^{+}}$. The dots with error bars are data combined from all energy points, and the red solid lines are the weighted PHSP MC corrected by the results of global fit. The blue dashed lines represent the distributions of the simulated events evenly distributed in PHSP, without polarization.
Figure 4 shows the distribution of the moment given by
$M(\cos\theta_{\Sigma^{+}})=\frac{m}{N}\sum_{i=1}^{N_{k}}(n^{i}_{1,y}-n^{i}_{2,y}),$ (7)
which is calculated for $m=10$ intervals in $\cos\theta_{\Sigma^{+}}$. Here, $N$ represents the total number of events in the data sample. According to ref. BESIII:2022yprl , $M(\cos\theta_{\Sigma^{+}})$ is related to the polarization by
$M(\cos\theta_{\Sigma^{+}})=\frac{\alpha_{\Sigma^{+}}-\alpha_{\bar{\Sigma}^{-}}}{2}\frac{1+\eta\cos^{2}\theta_{\Sigma^{+}}}{3+\eta}P_{y}(\theta_{\Sigma^{+}})$ (8)
Assuming $CP$ conservation, we have $\alpha_{\Sigma^{+}}=-\alpha_{\bar{\Sigma}^{-}}$, and the expected angular dependence of $M(\cos\theta_{\Sigma^{+}})$ is proportional to $\sqrt{1-\eta^{2}}\alpha_{\Sigma^{+}}\!\sin\Delta\Phi\cos\theta_{\Sigma^{+}}\sin\theta_{\Sigma^{+}}/(3+\eta)$, which is consistent with the data in figure 4.
Figure 4: The moment $M(\cos\theta_{\Sigma^{+}})$ as a function of $\cos\theta_{\Sigma^{+}}$ for the $e^{+}e^{-}\rightarrow\Sigma^{+}\bar{\Sigma}^{-}$ reaction at different CM energies. Points with error bars are data, the red solid lines are the weighted PHSP MC corrected by the results of global fit, and the blue dashed lines represent the distributions without polarization from simulated events, evenly distributed in the PHSP.
[…]
3.6820 404.0 $134$. $2^{+12.4}_{-11.8}$ $0.54\pm 0.17\pm 0.12$ $0.38\pm 0.40\pm 0.12$ $0.84\pm 0.20\pm 0.14$ 0.4
3.6830 28.7 $27$. $2^{+6.0}_{-5.4}$ $0.96\pm 0.13\pm 0.12$ $2.35\pm 1.66\pm 0.12$ $0.22\pm 0.37\pm 0.34$ 0.0
3.6840 28.7 $97$. $8^{+10.7}_{-10.1}$ $0.86\pm 0.15\pm 0.12$ $1.19\pm 0.61\pm 0.12$ $0.42\pm 0.21\pm 0.20$ 1.9
3.6850 26.0 $256$. $2^{+16.8}_{-16.2}$ $0.76\pm 0.10\pm 0.12$ $0.16\pm 0.32\pm 0.12$ $0.57\pm 0.15\pm 0.16$ 0.7
3.6870 25.1 $283$. $0^{+17.8}_{-16.8}$ $0.66\pm 0.12\pm 0.12$ $0.02\pm 0.27\pm 0.12$ $0.70\pm 0.15\pm 0.15$ 0.4
3.6910 69.4 $77$. $5^{+9.8}_{-8.8}$ $0.16\pm 0.23\pm 0.12$ $1.29\pm 0.54\pm 0.12$ $1.31\pm 0.30\pm 0.16$ 2.1
3.7100 70.3 $22$. $0^{+5.8}_{-4.7}$ $0.01\pm 0.40\pm 0.12$ $-2.64\pm 0.60\pm 0.12$ $1.55\pm 0.62\pm 0.19$ 1.4
AVG. - -- $0.69\pm 0.05\pm 0.05$ $0.14\pm 0.16\pm 0.05$ $0.72\pm 0.08\pm 0.07$-
5 Systematic uncertainty
The systematic uncertainties on the measurement of the $\Sigma^{+}$ hyperon polarization arise due to the $\Sigma^{+}(\bar{\Sigma}^{-})$ reconstruction, the requirements on the $p\pi^{0}(\bar{p}\pi^{0})$ mass window, the background estimation, the fit method, and the decay parameters $\alpha_{\Sigma^{+}/\bar{\Sigma}^{-}}$.
5.1 $\Sigma^{+}(\bar{\Sigma}^{-})$ reconstruction
The systematic uncertainty due to the $\Sigma^{+}(\bar{\Sigma}^{-})$ reconstruction efficiency, incorporating the efficiencies of tracking, PID, kinematic fit and requirements of $M_{\pi^{0}\pi^{0}}^{\rm Recoil}$ and $\pi^{0}$ reconstruction, is estimated using the control sample of $\psi(3686)\rightarrow\Sigma^{+}\bar{\Sigma}^{-}$ with the same method as described in refs. BESIII:2021gca ; BESIII:2022mfx ; BESIII:2024jgy . The efficiency difference is evaluated by comparing a full reconstruction sample that reconstructs both the baryon and anti-baryon sides, with two partial reconstruction samples that only reconstruct one side. The MC samples are then corrected with the efficiency difference, and the fit is repeated. The differences between the new and nominal values are taken as the systematic uncertainties.
5.2 Mass window
The uncertainty due to the requirements on the $p\pi^{0}(\bar{p}\pi^{0})$ mass window is estimated with the same method as introduced in ref. BESIII:2020fqg . The range of the $p\pi^{0}(\bar{p}\pi^{0})$ mass window is increased or decreased by 1 MeV/$c^{2}$, and the maximum differences between the new and nominal values are taken as the systematic uncertainties.
5.3 Background
The systematic uncertainty associated with the background is determined by comparing the results obtained from the fits with and without considering the sideband background. The modified likelihood function is expressed as
$\mathscr{S}=-\mathrm{ln}\mathscr{L}_{\rm data}+\mathrm{ln}\mathscr{L}_{\rm bg},$ (9)
where $\mathscr{L}_{\rm bg}$ is the likelihood function of background events determined in the sideband regions.
5.4 Fit method
To validate the reliability of the fit results, an input and output check is conducted using 500 pseudo-experiments. The helicity amplitude formula provided in ref. Faldt:2017kgy is utilized. To ensure that the Gaussian function is fitted comprehensively, the input values of the polarization parameters are deliberately chosen to be far away from the fit boundary. In order to ensure sufficient statistics, 6000 events are generated for each MC sample. In addition, to avoid the reconstruction effect, the MC samples used here have not been corrected by the detector. The differences between the input and output values obtained from the fits are negligible.
5.5 Decay parameter
The uncertainties from the decay parameters $\alpha_{\Sigma^{+}}$ for $\Sigma^{+}\to p\pi^{0}$ and $\alpha_{\bar{\Sigma}^{-}}$ for $\bar{\Sigma}^{-}\to\bar{p}\pi^{0}$ are estimated by varying the value, obtained from averaging the results in ref. BESIII:2020fqg , by ±$\pm 1\sigma$. The largest differences in the result are taken as the systematic uncertainties.
5.6 Total systematic uncertainty
The systematic uncertainties on the polarization measurement are summarized in table 2. Assuming all sources are independent, the total systematic uncertainties are determined by adding these sources in quadrature.
Table 2: Absolute systematic uncertainties of the measured parameters.
Source $\eta$ $\Delta\Phi$ (rad)
$\Sigma^{+}(\bar{\Sigma}^{-})$ reconstruction 0.02 0.00
Mass window 0.01 0.02
Sideband background 0.11 0.11
Fit method 0.00 0.00
Decay parameter 0.05 0.03
Total 0.12 0.12
6 Summary
In summary, based on $e^{+}e^{-}$ collision data corresponding to an integrated luminosity of 652.1 pb-1 collected with the BESIII detector at the BEPCII collider, we present an energy-dependent measurement of transverse polarization, relative phase and the modulus ratio of the form factors of $\Sigma^{+}$ hyperon in the $e^{+}e^{-}\to\Sigma^{+}\bar{\Sigma}^{-}$ reaction. For the first time, the phase of the $\Sigma^{+}$ hyperon electromagnetic form factors is explored in a higher range of four-momentum transfer above $q^{2}>13.5~{}\rm{GeV}^{2}$. No polarization is evident at each energy point. The comparison between this work and previous measurements BESIII:2023sgt ; BESIII:2023ynq is shown in figure 5. $\Delta\Phi$ is less than zero for $J/\psi$ decay and $\sqrt{s}=3.71$ GeV, and greater than zero at other energy points, which implies that there may be at least one $\Delta\Phi=0$ rad between these energy points. Such an evolution is important input for understanding its asymptotic behavior and the dynamics of baryons similar to the $\Lambda$ hyperon Mangoni:2021qmd . These results are important to understand the production mechanism of the $\Sigma^{+}$-$\bar{\Sigma}^{-}$ pairs at different energy points.
Figure 5: The comparisons of $\eta$, $\Delta\Phi$, and $R$ as a function of CM energy between this work and the previous measurements BESIII:2023sgt ; BESIII:2023ynq .
Acknowledgements.
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2020YFA0406400, 2020YFA0406300, 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 12075107, 12247101, 11635010, 11735014, 11835012, 11935015, 11935016, 11935018, 11961141012, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017; the 111 Project under Grant No. B20063; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; CAS Key Research Program of Frontier Sciences under Contracts Nos. QYZDJ-SSW-SLH003, QYZDJ-SSW-SLH040; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; European Union’s Horizon 2020 research and innovation programme under Marie Sklodowska-Curie grant agreement under Contract No. 894790; German Research Foundation DFG under Contracts Nos. 455635585, Collaborative Research Center CRC 1044, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374